Please help me. I am a two digit number. The sum of my digits is 9. The digit in the tens place is one half the value of the digit in the ones place. What number am I?
Answer by lwsshak3(11628)

I am a two digit number. The sum of my digits is 9. The digit in the tens place is one half the value of the digit in the ones place. What number am I.
***
let u=unit digit
t=tens digit
u+t=9
u=9-t
..
t=u/2
t=(9-t)/2
2t=9-t
3t=9
t=3
u=9-t=6
number: 36